# Dark Matter Substructure or Source Model Systematics? A Case Study of Cluster Lens Abell S1063

### v1

@@ Abstract
Mapping the small-scale structure of the universe through gravitational lensing is a promising tool for probing the particle nature of dark matter. Curved Arc Basis (CAB) has been proposed as a local lensing formalism in galaxy clusters, with the potential to detect low-mass dark matter substructure. In this work, we analyze the cluster lens Abell S1063 in search of dark matter substructure with the CAB formalism, using multi-band imaging data from JWST. We use two different source modeling methods: shapelets and pixel-based source reconstruction based on Delaunay triangulation. We find that source modeling systematics from shapelets result in a disagreement between Curved Arc Basis parameters measured from different filters. Source modeling with Delaunay significantly alleviates this systematic, as seen in the improvement in agreement across filters. We also find that inadequate complexity in source modeling can result in convincing spurious detections of dark matter substructure from strong gravitational lenses, as seen by our measurement of a subhalo with shapelets, a spurious detection that is not reproduced with Delaunay source modeling. We demonstrate that multi-band analysis with different JWST filters is key for disentangling source and lens model systematics from dark matter substructure detections.
@@
The standard model of cosmology (CDM) assumes the existence of Cold Dark Matter (CDM), which dominates the matter content of the universe and is essential to explaining an array of cosmological observables spanning cosmic epochs. These include observations of the cosmic microwave background (Planck Collaboration et al. 2020) and large-scale structure of the universe (Springel et al. 2006,Gil-Marín et al. 2016, Abbott et al. 2018, Troxel et al. 2018, Abolfathi et al. 2018, DESI Collaboration et al. 2024). While it has been remarkably successful at large scales, the cold dark matter paradigm has yet to be tested on small scales. Simulations with CDM predict an abundance of low-mass structures called subhalos residing in dark matter halos around galaxies (de Lucia et al. 2004, Chan et al. 2015). At the same time, many alternative dark matter theories can reproduce our large-scale observations but predict distinct observational signatures at sub-galactic scales. These theories either modify the predicted density profiles and population statistics of subhalos, as seen in warm or self-interacting dark matter models (Bode et al. 2001, Tulin & Yu 2018), or predict entirely different forms of substructure, such as interference patterns arising in ultralight axion dark matter (Schive et al. 2014). To this end, probing the small-scale structure of the universe can be the key to understanding the physics of dark matter.
@@
Strong gravitational lensing is sensitive to the line-of-sight mass distribution between the source and the observer (Mao & Schneider 1998), making it a direct probe of the dark matter substructure. As predicted by general relativity, light rays bend around massive objects, such as galaxies or galaxy clusters, leading to the appearance of multiple distorted images of a background source. The mass distribution of the lens is often parametrized by a smooth profile. A dark matter clump, either within the lens or along the line of sight, will cause a perturbation on top of this smooth lens model, allowing us to detect dark matter substructure gravitationally. To this day, a few subhalos have been detected or reported as candidate detections with this method in galaxy-galaxy strong lenses (Vegetti et al. 2010, Hezaveh et al. 2016, Şengül et al. 2022, Lange et al. 2024).
@@
Cluster lenses are even more powerful than galaxies, as they often host rich collections of multiply imaged, highly magnified, distorted arcs. However, studies of dark matter substructure in cluster lenses have been hindered by the computational cost of analyzing pixel-level data from their extensive fields of view. To this end, typical strong lens modeling of clusters has involved reducing the rich pixel-level data to only the positional information of lensed images (e.g., Lagattuta et al. 2019). However, recent work has made significant progress on incorporating pixel-level information in global cluster models (Acebron et al. 2024, Broadhurst et al. 2024). As a result, cluster substructure studies have been limited to statistical constraints on subhalo population abundances (Natarajan et al. 2007, Natarajan et al. 2017). While such studies give us important insights into the abundance of dark matter substructure, probing the density profiles of individual subhalos with direct detection techniques allows us to go a step further. By comparing the measured concentration, ellipticity, and effective density slope (Şengül & Dvorkin 2022) of subhalos to results from simulations with different dark matter models, we can support or rule out dark matter particle candidates and place tighter constraints on the viable parameter spaces of models.
@@
A promising approach for direct detection of cluster substructure is to employ a local lensing formalism. Curved Arc Basis (Birrer 2021) was proposed as a local description of gravitational distortions based on the eigenvalues and eigenvectors of the local lensing Jacobian. Şengül et al. (2023) showed with simulations that dark matter substructure along the line of sight of highly magnified lensed arcs in galaxy clusters can be detected and studied with the Curved Arc Basis (CAB) formalism. They also had the first application of CAB to real data by modeling lensed images of the first cluster lens SMACS 0723 observed by the James Webb Space Telescope (JWST).
@@
In this work, we exploit the power of multi-filter analysis of JWST data to study the source, lens, and foreground systematics involved in substructure searches with the Curved Arc Basis formalism. We describe the observations of Abell S1063 in Section 2. We detail our source and lens modeling methods, as well as substructure search in Section 3. We present our results of modeling a triply-lensed system in Abell S1063 with the Curved Arc Basis formalism and our substructure search in Section 4. We present our concluding thoughts in Section 5.
@@ 2.1 Triply-imaged system 4
We use the detailed studies of Abell S1063 to carefully select a set of multiple images best suited for study with the Curved Arc Basis formalism. We choose a set of three images identified as system 4 by Caminha et al. (2016) for a number of reasons:
@@ 2.2 JWST Observations
We use James Webb Space Telescope observations111https://www.stsci.edu/cgi-bin/get-proposal-info?id=1840&observatory=JWST of Abell S1063 proposed in the General Observing Program and taken on 24 September 2022 with the Near Infrared Camera (NIRCam). The data is available in three short-wavelength and three long-wavelength filters. We only analyze the short-wavelength filters (F115W, F150W, F200W) as they offer a sharper point spread function and twofold pixel resolution compared to the longer-wavelength filters. Both features are important to our analysis, as the substructure search relies on distinguishing pixel-level changes in the brightness of lensed images. The broadband filters F115W, F150W, and F200W are sensitive to a range of non-overlapping frequencies centered at pivot wavelengths of 1.154, 1.501, and 1.988 m, respectively. Thereby, each filter provides an independent realization of varying source morphologies lensed by the same mass distribution. Each filter’s exposure times and signal-to-noise ratios are reported in Table 1. The image cutouts of system 4 in the three filters are shown in Figure 2. The data used in our analysis is publicly available on the Mikulski Archive for Space Telescopes (MAST)222https://mast.stsci.edu/portal/Mashup/Clients/Mast/Portal.html.
@@
We use the simulation tool WebbPSF333https://www.stsci.edu/jwst/science-planning/proposal-planning-toolbox/psf-simulation-tool to model the point spread function (PSF) of each NIRCam filter. Typically, the PSF can also be calculated from pointlike sources in the field of view. However, we turn to simulation software because field stars are scarce along the line of sight of Abell S1063, located at a galactic latitude of , in a direction away from the galactic plane.
@@ 3 Methods
Strong lensing analysis requires simultaneous modeling of the source light and lens-mass distribution, obtained by minimizing residuals between the observed data and model reconstructions. Our data (see Figure 2) consist of elliptical cutouts from pixel images of system 4 in Abell S1063 taken in three different filters. To model the cluster lensing, we use the local lensing formalism of Curved Arc Basis (Birrer 2021), which has been shown to work well in describing highly magnified, extended arcs in cluster lenses, as well as to be sensitive to dark matter substructure (Şengül et al. 2023). To capture the light distribution of the background galaxy, we use two alternative source modeling methods: parametric shapelets (Refregier 2003, Birrer et al. 2015, Birrer et al. 2021) and pixel-based Delaunay triangulation (Vegetti & Koopmans 2009, Şengül 2024). To perform statistical fitting and search for dark matter substructure, we use Bayesian nested sampling with dynesty (Speagle 2020).
3.1 Lens Modeling: Curved Arc Basis
3.1.1 Angular Deflection in Curved Arc Basis
The Curved Arc Basis formalism proposed by Birrer (2021) uses the eigenvalues and eigenvectors of the local lensing Jacobian to describe distortions of extended arcs. This local model is fully equivalent to a global Singular Isothermal Sphere with a mass sheet transformation (MST)444 MST is a transformation of the form , where and is the convergence field. Such transformations lead to a mass-sheet degeneracy as the shear signal is invariant under the addition of a constant surface mass density to the convergence field. and can be described by four local parameters: tangential stretch , radial stretch , orientation , and curvature . An illustration of the effect of a Curved Arc Basis distortion on a circularly symmetric source is shown in Figure 3. The deflection angles of a Curved Arc Basis transformation centered at can be expressed as:
@@ 3.1.1 Angular Deflection in Curved Arc Basis
The Curved Arc Basis formalism is especially relevant in cluster lenses, where the most massive component of the cluster dominates the strong gravitational lensing effect. Yang et al. (2020) have shown that, in some cases, lensed images in clusters may be well described by a constant convergence and shear, which is a special limit of the Curved Arc Basis as . Şengül et al. (2023) first applied CAB to data, modeling JWST images of the cluster lens SMACS 0723, and presented a novel method for detecting and studying low-mass subhalos. In particular, with the local lensing formalism in hand, one can detect subhalos located near highly magnified arcs by discerning their perturbative corrections to the main lensing by the cluster. In mock JWST-quality simulations of the SMACS 0723 system, the authors showed that the CAB formalism could detect perturbers down to masses of , allowing a measurement of their concentration, ellipticity, and redshift (Şengül et al. 2023).
@@ 3.1.2 Region of Validity
Curved Arc Basis is a local formalism and is only able to perfectly capture the lensing distortions when the global model is a Singular Isothermal Sphere (SIS), as given in Equation (1). Lenses that are more complex than SIS will add corrections to CAB.
The ellipticity of the most massive cluster dark matter halo will induce derivatives along the tangential stretch of a curved arc. External shear will cause additional asymmetric distortions to the Curved Arc Basis lensing. These two effects can be accounted for by considering extensions of Curved Arc Basis incorporating tangential derivatives and shear, which add corrections to regular CAB away from the center of the image (Birrer 2021). However, the true lensing in a cluster is more complex than Curved Arc Basis and its extensions: the real mass distribution of clusters does not precisely follow these parametric profiles, an effect which can become important locally; at the same time, nearby galaxies in the field of view of the lensed images can induce additional lensing distortions to the images. The masses and concentrations of these galaxies are often poorly constrained, and incorporating them into the lens modeling can make the problem computationally intractable. Therefore, it has been emphasized that Curved Arc Basis is a local formalism that should be used cautiously (Birrer 2021). The region of local validity of a CAB approximation can easily be studied in simulations. However, the conclusions on the validity of CAB are heavily dependent on the specific configuration of our lensed images relative to the complex mass and light distribution of the cluster in question. In many cases, a detailed global model of a cluster lens may not be available. In this work, we take a general, empirical approach to illustrate the power of multi-band analysis with JWST in determining the region of CAB validity and disentangling lens model systematics. To this end, we perform our analysis in two different mask sizes applied to our JWST images. Masks are elliptical boundaries on the lens plane within which all the pixels are included in our analysis and outside of which all the pixels are excluded, illustrated with white dashed elliptical contours in Figure 2. If the local lensing formalism is a good approximation, assuming all other systematics are negligible, the parameters measured in different filters should be consistent. If the formalism fails due to inadequacies of CAB away from image centers, we should expect to see a tension between the same parameters measured in different filters, since different filters will have different spatial extents of source light. However, systematics can also arise from source modeling. To distinguish the two sources of systematics, we employ two different source modeling methods, as discussed in Section 3.2.
@@ 3.2 Source Modeling
Source modeling is an important part of strong gravitational lensing analysis, but it is especially important for Curved Arc Basis studies. JWST cluster data often involves more strongly magnified and higher-resolution lensed images than typical galaxy-galaxy strong lensing systems, thereby demanding a more sophisticated source model. In this work, we explore the strengths and limitations of two different source modeling methods: parametric shapelets and pixel-based Delaunay triangulation.
@@ 3.2.1 Shapelets
Shapelets (Refregier 2003, Birrer et al. 2015, Birrer et al. 2021) are an orthonormal set of weighted Hermite polynomials. A shapelet source model comprises a linear combination of the polynomial basis functions and is thereby fully specified by a list of shapelet coefficients, a reference scale, and a coordinate center. Shapelets are able to capture more complex light distributions than the common Sérsic (Sérsic 1963) profile of galaxies and have thus been commonly used in strong lensing analyses. The complexity of the source model is controlled by the shapelet basis order , which determines the number of allowed basis functions through the relation . The size of the source is set by the reference scale parameter . The smallest scale resolved by order and reference scale is given by the relation . Finally, the light peak position of the source model is controlled by the shapelet center , which is a non-linear parameter in the model.
We determine the posterior distributions of and through nested sampling together with the lens model parameters, using uniform priors. At each sampling step, where the lens model is fixed, the light distribution on the lens plane is a simple linear combination of the shapelet basis function coefficients on the source plane. Thus, we can solve for the shapelet coefficients by a linear matrix inversion of an matrix, where the length of the data vector is the total number of pixels in our image cutouts. As the shapelet basis order is a hyperparameter, we perform a Bayesian Information Criterion (BIC) analysis to find the optimal with the lowest , where is the number of model parameters. This allows us to ensure that we use the optimal source complexity without overfitting the noise in our data. We perform separate BIC analyses for the shapelet basis order for each NIRCam filter, as the source model should naturally look different in different wavelength ranges in each filter, as well as for each lens model and mask size applied to image cutouts. All reported results use the shapelet that gives model fits with the lowest BIC score.

### v2

@@ Abstract
Mapping the small-scale structure of the universe through gravitational lensing is a promising tool for probing the particle nature of dark matter. Curved Arc Basis (CAB) has been proposed as a local lensing formalism in galaxy clusters, with the potential to detect low-mass dark matter substructure. In this work, we analyze the cluster lens Abell S1063 in search of dark matter substructure with the CAB formalism, using multi-band imaging data from JWST. We use two different source modeling methods: shapelets and pixel-based source reconstruction based on Delaunay triangulation. We find that source modeling systematics from shapelets result in a disagreement between CAB parameters measured from different filters. Source modeling with Delaunay significantly alleviates this systematic, as seen in the improvement in agreement across filters. We also find that inadequate complexity in source modeling can result in convincing spurious detections of dark matter substructure from strong gravitational lenses, as seen by our measurement of a subhalo with shapelets, a spurious detection that is not reproduced with Delaunay source modeling. We demonstrate that multi-band analysis with different JWST filters is key for disentangling source and lens model systematics from dark matter substructure detections.
@@
The standard model of cosmology (CDM) assumes the existence of Cold Dark Matter (CDM), which dominates the matter content of the universe and is essential to explaining an array of cosmological observables spanning cosmic epochs. These include observations of the cosmic microwave background (Planck Collaboration et al. 2020) and large-scale structure of the universe (Springel et al. 2006,Gil-Marín et al. 2016, Abbott et al. 2018, Troxel et al. 2018, Abolfathi et al. 2018, Adame et al. 2025). While it has been remarkably successful at large scales, the cold dark matter paradigm has yet to be tested on small scales. Simulations with CDM predict an abundance of low-mass structures called subhalos residing in dark matter halos around galaxies (de Lucia et al. 2004, Chan et al. 2015). At the same time, many alternative dark matter theories can reproduce our large-scale observations but predict distinct observational signatures at sub-galactic scales. These theories either modify the predicted density profiles and population statistics of subhalos, as seen in warm or self-interacting dark matter models (Bode et al. 2001, Tulin & Yu 2018), or predict entirely different forms of substructure, such as interference patterns arising in ultralight axion dark matter (Schive et al. 2014). To this end, probing the small-scale structure of the universe can be the key to understanding the physics of dark matter.
@@
Strong gravitational lensing is sensitive to the line-of-sight mass distribution between the source and the observer (Mao & Schneider 1998), making it a direct probe of the dark matter substructure. As predicted by general relativity, light rays bend around massive objects, such as galaxies or galaxy clusters, leading to the appearance of multiple distorted images of a background source. The mass distribution of the lens is often parametrized by a smooth profile. A dark matter clump, either within the lens or along the line of sight, will cause a perturbation on top of this smooth lens model, allowing us to detect dark matter substructure gravitationally. To this day, a few subhalos have been detected or reported as candidate detections with this method in galaxy-galaxy strong lenses (Vegetti et al. 2010, Hezaveh et al. 2016, Şengül et al. 2022, Lange et al. 2025).
@@
Cluster lenses are even more powerful than galaxies, as they often host rich collections of multiply imaged, highly magnified, distorted arcs. However, studies of dark matter substructure in cluster lenses have been hindered by the computational cost of analyzing pixel-level data from their extensive fields of view. To this end, typical strong lens modeling of clusters has involved reducing the rich pixel-level data to only the positional information of lensed images (e.g., Lagattuta et al. 2019). However, recent work has made significant progress on incorporating pixel-level information in global cluster models (Acebron et al. 2024, Broadhurst et al. 2025). As a result, cluster substructure studies have been limited to statistical constraints on subhalo population abundances (Natarajan et al. 2007, Natarajan et al. 2017). While such studies give us important insights into the abundance of dark matter substructure, probing the density profiles of individual subhalos with direct detection techniques allows us to go a step further. By comparing the measured concentration, ellipticity, and effective density slope (Şengül & Dvorkin 2022) of subhalos to results from simulations with different dark matter models, we can support or rule out dark matter particle candidates and place tighter constraints on the viable parameter spaces of models.
@@
A promising approach for direct detection of cluster substructure is to employ a local lensing formalism. Curved Arc Basis (CAB) (Birrer 2021) was proposed as a local description of gravitational distortions based on the eigenvalues and eigenvectors of the local lensing Jacobian, the curvature of the tangential eigenvector, and an orientation of the curved arc. Şengül et al. (2023) showed with simulations that dark matter substructure along the line of sight of highly magnified lensed arcs in galaxy clusters can be detected and studied with the CAB formalism. They also had the first application of CAB to real data by modeling lensed images of the first cluster lens SMACS 0723 observed by the James Webb Space Telescope (JWST).
@@
In this work, we exploit the power of multi-filter analysis of JWST data to study the source, lens, and foreground systematics involved in substructure searches with the CAB formalism. We describe the observations of Abell S1063 in Section 2. We detail our source and lens modeling methods, as well as substructure search in Section 3. We present our results of modeling a triply-lensed system in Abell S1063 with the CAB formalism and our substructure search in Section 4. We present our concluding thoughts in Section 5.
@@ 2.1 Triply-imaged system 4
We use the detailed studies of Abell S1063 to carefully select a set of multiple images best suited for study with the CAB formalism. We choose a set of three images identified as system 4 by Caminha et al. (2016) for a number of reasons:
@@ 2.2 JWST Observations
We use James Webb Space Telescope observations111http://dx.doi.org/10.17909/mskr-6n50 of Abell S1063 proposed in the General Observing Program under Program 222https://www.stsci.edu/jwst/science-execution/program-information?id=1840 and taken on 24 September 2022 with the Near Infrared Camera (NIRCam). The data is available in three short-wavelength and three long-wavelength filters. We only analyze the short-wavelength filters (F115W, F150W, F200W) as they offer a sharper point spread function and twofold pixel resolution compared to the longer-wavelength filters. Both features are important to our analysis, as the substructure search relies on distinguishing pixel-level changes in the brightness of lensed images. The broadband filters F115W, F150W, and F200W are sensitive to a range of non-overlapping frequencies centered at pivot wavelengths of 1.154, 1.501, and 1.988 m, respectively. Thereby, each filter provides correlated but varying source morphologies lensed by the same mass distribution. Each filter’s exposure times and signal-to-noise ratios are reported in Table 1. The image cutouts of system 4 in the three filters are shown in Figure 2. The data used in our analysis is publicly available on the Mikulski Archive for Space Telescopes (MAST)333https://mast.stsci.edu/portal/Mashup/Clients/Mast/Portal.html.
@@
We use the simulation tool WebbPSF444https://www.stsci.edu/jwst/science-planning/proposal-planning-toolbox/psf-simulation-tool to model the point spread function (PSF) of each NIRCam filter. Typically, the PSF can also be calculated from pointlike sources in the field of view. However, we turn to simulation software because field stars are scarce along the line of sight of Abell S1063, located at a galactic latitude of , in a direction away from the galactic plane.
@@ 3 Methods
Strong lensing analysis requires simultaneous modeling of the source light and lens-mass distribution, obtained by minimizing residuals between the observed data and model reconstructions. Our data (see Figure 2) consist of elliptical cutouts from pixel images of system 4 in Abell S1063 taken in three different filters. To model the cluster lensing, we use the local lensing formalism of CAB (Birrer 2021), which has been shown to work well in describing highly magnified, extended arcs in cluster lenses, as well as to be sensitive to dark matter substructure (Şengül et al. 2023). To capture the light distribution of the background galaxy, we use two alternative source modeling methods: parametric shapelets (Refregier 2003, Birrer et al. 2015, Birrer et al. 2021) and pixel-based Delaunay triangulation (Vegetti & Koopmans 2009, Şengül 2025). To perform statistical fitting and search for dark matter substructure, we use Bayesian nested sampling with dynesty (Speagle 2020).
3.1 Lens Modeling: CAB
3.1.1 Angular Deflection in CAB
The CAB formalism proposed by Birrer (2021) uses the eigenvalues and eigenvectors of the local lensing Jacobian to describe distortions of extended arcs. This local model is fully equivalent to a global Singular Isothermal Sphere with a mass sheet transformation (MST)555 MST is a transformation of the form , where and is the convergence field. Such transformations lead to a mass-sheet degeneracy as the shear signal is invariant under the addition of a constant surface mass density to the convergence field. and can be described by four local parameters: tangential stretch , radial stretch , orientation , and curvature . An illustration of the effect of a CAB distortion on a circularly symmetric source is shown in Figure 3. The deflection angles of a CAB transformation centered at can be expressed as:
@@ 3.1.1 Angular Deflection in CAB
The CAB formalism is especially relevant in cluster lenses, where the most massive component of the cluster dominates the strong gravitational lensing effect. Yang et al. (2020) have shown that, in some cases, lensed images in clusters may be well described by a constant convergence and shear, which is a special limit of the CAB as . Şengül et al. (2023) first applied CAB to data, modeling JWST images of the cluster lens SMACS 0723, and presented a novel method for detecting and studying low-mass subhalos. In particular, with the local lensing formalism in hand, one can detect subhalos located near highly magnified arcs by discerning their perturbative corrections to the main lensing by the cluster. In mock JWST-quality simulations of the SMACS 0723 system, the authors showed that the CAB formalism could detect perturbers down to masses of , allowing a measurement of their concentration, ellipticity, and redshift (Şengül et al. 2023).
@@ 3.1.2 Region of Validity
CAB is a local formalism and is only able to perfectly capture the lensing distortions for a global model that consists of a Singular Isothermal Sphere (SIS) and an allowed Mass Sheet Transform (MST), as given in Equation (1). Lenses that are more complex than SIS and MST will add corrections to CAB.
The ellipticity of the most massive cluster dark matter halo will induce derivatives along the tangential stretch of a curved arc. External shear will cause additional asymmetric distortions to the CAB lensing. These two effects can be accounted for by considering extensions of CAB incorporating tangential derivatives and shear, which add corrections to regular CAB away from the center of the image (Birrer 2021). However, the true lensing in a cluster is more complex than CAB and its extensions: the real mass distribution of clusters does not precisely follow these parametric profiles, an effect which can become important locally; at the same time, nearby galaxies in the field of view of the lensed images can induce additional lensing distortions to the images. The masses and concentrations of these galaxies are often poorly constrained, and incorporating them into the lens modeling can make the problem computationally intractable. Therefore, it has been emphasized that CAB is a local formalism that should be used cautiously (Birrer 2021). The region of local validity of a CAB approximation can easily be studied in simulations. However, the conclusions on the validity of CAB are heavily dependent on the specific configuration of our lensed images relative to the complex mass and light distribution of the cluster in question. In many cases, a detailed global model of a cluster lens may not be available. In this work, we take a general, empirical approach to illustrate the power of multi-band analysis with JWST in determining the region of CAB validity and disentangling lens model systematics. To this end, we perform our analysis in two different mask sizes applied to our JWST images. Masks are elliptical boundaries on the lens plane within which all the pixels are included in our analysis and outside of which all the pixels are excluded, illustrated with white dashed elliptical contours in Figure 2. If the local lensing formalism is a good approximation, assuming all other systematics are negligible, the parameters measured in different filters should be consistent. If the formalism fails due to inadequacies of CAB away from image centers, we should expect to see a tension between the same parameters measured in different filters, since different filters will have different spatial extents of source light. However, systematics can also arise from source modeling. To distinguish the two sources of systematics, we employ two different source modeling methods, as discussed in Section 3.2.
@@ 3.2 Source Modeling
Source modeling is an important part of strong gravitational lensing analysis, but it is especially important for CAB studies. JWST cluster data often involves more strongly magnified and higher-resolution lensed images than typical galaxy-galaxy strong lensing systems, thereby demanding a more sophisticated source model. In this work, we explore the strengths and limitations of two different source modeling methods: parametric shapelets and pixel-based Delaunay triangulation.
@@ 3.2.1 Shapelets
Shapelets (Refregier 2003, Birrer et al. 2015, Birrer et al. 2021) are an orthonormal set of weighted Hermite polynomials. A shapelet source model comprises a linear combination of the polynomial basis functions and is thereby fully specified by a list of shapelet coefficients, a reference scale, and a coordinate center. Shapelets are able to capture more complex light distributions than the common Sérsic (Sérsic 1963) profile of galaxies and have thus been commonmber of allowed basis functions through the relation . The size of the source is set by the reference scale parameter . The smallest scale resolved by order and reference scale is given by the relation . Finally, the light peak position of the source model is controlled by the shapelet center , which is a non-linear parameter in the model.
We determine the posterior distributions of and through nested sampling together with the lens model parameters, using uniform priors. At each sampling step, where the lens model is fixed, the light distribution on the lens plane is a simple linear combination of the shapelet basis function coefficients on the source plane. Thus, we can solve for the shapelet coefficients by a linear matrix inversion of an matrix, where the length of the data vector is the total number of pixels in our image cutouts. As the basis order is a hyperparameter, we perform a Bayesian Information Criterion (BIC) analysis to find the optimal with the lowest , where is the number of model parameters. This allows us to ensure that we use the optimal source complexity without overate BIC analyses for the shapelet basis order for each NIRCam filter, as the source model should naturally look different in different wavelength ranges in each filter, as well as for each lens model and mask size applied to image cutouts. All reported results use the shapelet that gives model fits with the lowest BIC score.
In our fiducial analysis with shapelets, we use a Cartesian basis, which is commonly used in lens modeling but is better suited to spherical source morphologies. However, if the source is highly elliptical, the implicit prior set by our choice of basis can bias the results. For this reason, we repeat our analysis using source reconstruction with an elliptical shapelet basis, which introduces two additional non-linear parameters. We set uninformative priors on these two parameters, allowing source ellipticity and position-angle of the semi-major axis to vary between and . Henceforth, refer to these two shapelet source reconstruction schemes with different bases as ‘Cartesian shapelets’ and ‘elliptical shapelets’. When a statement applies to both methods, we refer to them collectively as ‘shapelets’.